# Efficient Multi-Agent Collaboration with Tool Use for Online Planning in Complex Table Question Answering

### v1

@@ Abstract
Complex table question answering (TQA) aims to answer questions that require complex reasoning, such as multi-step or multi-category reasoning, over data represented in tabular form. Previous approaches demonstrated notable performance by leveraging either closed-source large language models (LLMs) or fine-tuned open-weight LLMs. However, fine-tuning LLMs requires high-quality training data, which is costly to obtain, and utilizing closed-source LLMs poses accessibility challenges and leads to reproducibility issues. In this paper, we propose Multi-Agent Collaboration with Tool use (MACT), a framework that requires neither closed-source models nor fine-tuning. In MACT, a planning agent and a coding agent that also make use of tools collaborate to answer questions. Our experiments on four TQA benchmarks show that MACT outperforms previous SoTA systems on three out of four benchmarks and that it performs comparably to the larger and more expensive closed-source model GPT-4 on two benchmarks, even when using only open-weight models without any fine-tuning. We conduct extensive analyses to prove the effectiveness of MACT’s multi-agent collaboration in TQA.
@@ 1 Introduction
The goal of table question answering (TQA) is to answer a question based on data represented in tabular form, optionally also using additional textual context. Recent studies on TQA focus more and more on complex instances, as they are ubiquitous in table data analysis (Zhu et al., 2021; Zhang et al., 2024b; Lu et al., 2023). Solving those complex instances requires performing multiple reasoning steps and/or employing different reasoning strategies (Ghosal et al., 2023). We refer to these aspects as “multi-step” and “multi-category” reasoning, respectively. An example requiring both multi-step and multi-category reasoning is shown in the upper left part of Figure 1. To answer the question, a system needs to first use factual knowledge to extract countries in Europe. Then, numerical reasoning is applied to calculate the percentage change and carry out the comparison.
@@
One popular approach for addressing those complex instances in TQA is planning, where step-wise plans are generated and used to guide the reasoning process (Zhang et al., 2024c; Wang et al., 2024; Wu and Feng, 2024; Zhu et al., 2024; Zhao et al., 2024). State-of-the-art works in this direction either fine-tune open-weight large language models (LLMs) Wu and Feng (2024); Zhu et al. (2024) or prompt closed-source commercial LLMs Wang et al. (2024); Zhang et al. (2024c); Zhao et al. (2024). However, fine-tuning requires high-quality data, which is usually hard to obtain, while prompting closed-source commercial LLMs can be costly and poses challenges to reproducibility. Furthermore, these methods rely on a single LLM, hence they struggle with performing different types of reasoning. To the best of our knowledge, existing models leverage a single LLM to perform planning and reasoning, which is sub-optimal in particular if the LLM does not excel at mathematical reasoning or coding (Wu and Feng, 2024).
To address these challenges, we propose a multi-agent collaboration framework with tool use. Our framework neither depends on closed-source LLMs nor requires fine-tuning. In fact, its backbone LLMs can be exchanged flexibly. It incorporates two agents (a planning agent and a coding agent) and a set of tools (a Python interpreter, a calculator and Wikipedia search). The planning agent performs online planning, i.e., it generates a plan iteratively. This breaks down complex problems and helps address multi-step reasoning. The coding agent and the tool set assist with generating faithful intermediate results. We let these agents work in a collaborative setting, addressing the challenges of multi-category reasoning as all agents can concentrate on the reasoning types they excel in. We further introduce an efficiency optimization module which allows the framework to take shortcuts. We conduct experiments on four popular TQA benchmarks that include complex TQA instances. Our framework outperforms previous SoTA systems on three out of four benchmarks. It achieves comparable results to GPT-4 on two benchmarks even when using only open-weight models without any fine-tuning. In comparison to fine-tuned SoTA TQA systems, it demonstrates considerably better generalizability. Our analysis proves the effectiveness of our proposed collaborative setting of specialized agents. We further show that the efficiency optimization module can save up to 33% of iterations without performance degradation.
@@ 2 Related Work
Table 1: Comparing MACT with previous works. NL plan stands for using natural language to encode a plan or not. cal and Wiki refer to calculator and Wikipedia.
@@
We categorize previous work into three groups based on planning strategies: Heuristic coarse-grained planning consists of the same two steps retrieving and aggregation for all questions Ye et al. (2023); Zhou et al. (2024). Online global planning generates a plan in the first iteration and revises it in the next one Zhao et al. (2024). Online iterative planning conditions the generation of the next step on the executed results of previous steps Zhang et al. (2023a); Wang et al. (2024). We opt for online iterative planning for complex TQA, as it introduces more granularities (in contrast to heuristic planning) and emphasizes the dependency among steps (in contrast to online global planning). Focusing on the dependency among steps allows a system to look into a problem in depth. This aligns with the task profile of complex TQA. In contrast to previous work using iterative planning, we introduce an efficiency optimization module to minimize the costs of the framework. To learn plan generation, previous work either depends on fine-tuning (Wu and Feng, 2024; Zhu et al., 2024), or strong closed-source models, combined with in-context learning (Wang et al., 2024; Zhang et al., 2023a; Zhao et al., 2024). By contrast, MACT is suitable for both closed-source and open-weight models, without the need for fine-tuning.
@@
In multi-agent collaboration settings, multiple AI entities collaborate towards a common goal (Talebirad and Nadiri, 2023). We use the term agent to refer to LLMs that can interact with executable tools, following Qiao et al. (2024). As far as we know, none of the previous works in TQA utilize multi-agent collaboration. The most related works explore collaboration among the same set of agents, i.e., all agents use the same backbone but are prompted differently (Liu et al., 2023; Zhao et al., 2024). The effectiveness of this approach relies heavily on the strong backbone models, which usually are closed-sourced. Our work explores multi-agent collaboration for TQA, without any constraint of model type.
@@ 3 Method
We propose MACT, a Multi-Agent Collaboration framework enriched with a set of Tools for TQA. Figure 1 provides an overview of the framework. It consists of four major modules: a memory $S$, a planning agent $M_{p}$, a coding agent $M_{c}$, and a tool set $T$. $M_{p}$ and $M_{c}$ are instantiated by (potentially) different LLMs. They collaborate through five core stages: action generation, action selection, tool selection/code creation, observation computation, and memory state update. The stages are executed iteratively for a maximum of $I$ iterations, where $I$ is a hyper-parameter. We control the overall efficiency of the collaboration via an efficiency optimization module. For a new TQA instance, we initialize the memory state $s_{0}=(\text{table},\text{question},\text{texts})$, i.e., with the input table, the question, and (if given) textual context. [^1] Some TQA datasets, such as TAT, provide textual context in addition to tables and questions. All parts of the memory are represented as strings. To represent the table as a string, we use pipes as column separators.
@@ 3.1 Action Generation
Retrieval textual description of what to retrieve, e.g., “sell numbers of 2019” $t=Python$ $M_{c}$ Python interpreter is run on generated code
@@
The intents Retrieval and Calculation are commonly seen in previous works (Gemmell and Dalton, 2023; Zhu et al., 2024). We use Retrieval for any operations extracting information from a table, including direct querying, filtering, and grouping. Instructions that require calculation, counting or comparison are captured by Calculation. To fulfill the possible need for external (factual) knowledge that is not present in the table or textual context, we add the intent Search. Read covers the need for contextual reasoning in table-text QA. It refers to instructions involving retrieving information from the texts provided as part of TQA instances. The intent Finish stops $M_{p}$ from generating more actions and ends the iterative execution of our framework, providing the final answer in the corresponding instruction. Lastly, we use an intent called Ask to retrieve an answer based on the internal knowledge from the planning agent.
@@ 3.3 Tool Selection and Use
The tool needed for executing action $a^{*}_{i}$ depends on the intent of the action (see columns “Tool Selection” and “Tool Use/Execution” of Table 2). To address the intents Search, Calculation and Retrieval, we introduce a set of tools $T=\{T_{srh},T_{cal}\,T_{python}\}$. $T_{srh}$ is an API function for Wikipedia search (Shinn et al., 2023) from Langchain. [^2] https://rubydoc.info/gems/langchainrb/Langchain/Tool/Wikipedia The API takes a target entity specified in the instruction and returns the first paragraph of the corresponding Wikipedia entry. $T_{cal}$ is a calculator, powered by a Python interpreter. It takes a formula generated by $M_{p}$ and outputs the answer. Note that the instruction of Calculation can also be a textual description, such as “Compute the average number of medals for each country in the table.” To better address these instructions, we introduce a coding agent $M_{c}$ with a Python interpreter, denoted as $T_{python}$ (see 3.4). The intent Retrieval is also addressed by $M_{c}$ and $T_{python}$, i.e., $M_{c}$ generates Python code based on a given instruction to retrieve target cells in the tables and the Python interpreter returns the executed results. Lastly, For Read, Ask and Finish, no tool is used, denoted as $t=Null$ in Table 2. The answers to the intents Read and Ask are queried via $M_{p}$. For Read, $M_{p}$ reads the answer from a given textual context. For Ask, it responds based on its internal knowledge. No execution is performed for Finish, as the intent ends MACT with a final answer in its instruction.
@@ 4 Experiments
We categorize the recent TQA systems into two groups indicating if they require LLM fine-tuning or not. As systems requiring fine-tuning, we include OmniTab (Jiang et al., 2022), TableLlama Zhang et al. (2024a), ProTrix (Wu and Feng, 2024), TAT-LLM (Zhu et al., 2024) and TableLLM Zhang et al. (2024b). Except for OmniTab, which is backboned by BART (Lewis et al., 2020), all others build on top of LLaMA 7b (Touvron et al., 2023). As TQA systems that do not require fine-tuning, we include Dater (Ye et al., 2023), Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2024c), TabSQLify (Nahid and Rafiei, 2024), Plan-then-Reason (Wu and Feng, 2024), Mix-SC (Liu et al., 2023) and ARC Zhang et al. (2023b). They all use GPT-3.5-turbo.
@@
In MACT, the choice of the planning and coding agent is flexible as no fine-tuning is involved. We experiment with the best open-weight LLMs available at the time of writing: Qwen 72B (Yang et al., 2024) (planning agent) and CodeLLaMA-34B (coding agent). They are run on int4 and full precision, respectively. We further use GPT-3.5-turbo as both the planning and coding agents when comparing our method with other TQA systems that use this model. $\tau_{p}$ and $\tau_{c}$ are set to 0.6 for non-repetitive action generation. We set the action and code generation size $k$ to 5, following Liu et al. (2023). The maximum number of iteration $I$ is set to 7, based on empirical results on the development sets. For the efficiency component, we set $\alpha$ to 1 to ensure high confidence of the model. We explore the effects of different values of $\alpha$ in Section 6. Two NVIDIA A100 GPUs are used for running MACT.
@@ 5 Results
We first evaluate the performance of MACT in direct comparison to recent TQA systems using closed-source LLMs. Second, we examine how our method performs compared to fine-tuned open-weight LLMs. Following prior work, we use exact match (EM) as the evaluation measure. Lastly, we discuss the efficiency of MACT.
@@
Table 3: Results of models without fine-tuning. SCT refers to SCITAB. The models are grouped by the LLM they use as their backbone (top: GPT-3.5; bottom: open-weight LLMs as indicated in parentheses). Performances marked with * are taken from the original paper. Performances marked with †are taken from Wu and Feng (2024), Zhu et al. (2024), Zhang et al. (2023b) and Lu et al. (2023) for each dataset.
@@
The middle part of Table 3 provides the results of MACT using specialized agents (MACT (Qw+CL): Qwen as planning agent, CodeLLaMA as coding agent). As baselines, we compare a setting without a specialized coding agent (MACT (Qw+Qw)) and a setting without a general planning agent (MACT(CL+CL)). As further baselines, we use the two LLMs on their own as well in combination, with Chain-of-thought prompting (Wei et al., 2022). For combination, the single models are prompted five times and combined using SC, as in Liu et al. (2023). This is a direct multi-agent baseline without collaboration and tool use. Both SC(Qw+CL) and MACT (Qw+CL) achieve higher EM scores than individually prompting Qwen and CodeLLaMA, demonstrating the positive effect of using multiple agents for planning and coding. Importantly, MACT (Qw+CL) outperforms SC(Qw+CL) by approximately 6 EM points on average across all datasets, highlighting the superiority of our collaboration technique over simply taking the most frequent predictions from two independent agents. We also find that having an expert coding agent for code generation (MACT (Qw+Qw) vs. MACT (Qw+CL)) improves performance considerably.
@@
TableLlama (LlaMA 7b) c 29.9 17.4 26.9 38.6
Protrix (LlaMA 7b) c 48.9* 26.8 40.2 42.4
TAT-LLM (LlaMA 7b) c - 69.6* - -
@@
Table 4: Results of MACT using different LLM agents in comparison to fine-tuned TQA models. Performances marked with * refer to the in-domain setting (where fine-tuning took place). SCT refers to SCITAB. CS refers to the deepseek-coder model.
MACT with open-source models delivers comparable performances as closed-source systems.
As shown in Table 3, by comparing MACT (Qw+CL) with TQA systems that rely on closed-source LLMs (in the upper part of Table 3), we find that our model outperforms the examined TQA systems for three out of four datasets. This is an important finding, as TQA systems based on commercial closed-source LLMs are costly to operate and often lack reproducibility. In contrast, our multi-agent TQA system is more cost-efficient and straightforward to replicate, while delivering superior performance compared to closed-source systems. In addition, we show two upper-bounds in the bottom part of Table 3: The performance of human annotators and of GPT-4. We observe that MACT achieves comparable performance as GPT-4 on two datasets, without the need for fine-tuning.
Table 4 compares our framework with prior fine-tuned TQA models. We present results for MACT with different planning and coding agents: Our standard setting (LlaMA-7b+CS-7b) as well as with a stronger planner (Qw-7b+CS-7b). For the models from related work, we mark the dataset on which they have been fine-tuned with * signs. In general, their performance on the dataset used for fine-tuning is rather high while they suffer from a considerable drop in EM when tested on other datasets. This observation is in line with the findings by (Zhang et al., 2024a; Huang et al., 2024). In contrast, MACT does not use fine-tuned models and can, thus, be applied to any dataset with a good generalization performance. MACT demonstrates comparable results as Protrix when using LlaMA-7b as the planning agent, though it has not been fine-tuned. As expected, using a better planning agent leads to better results. This also shows the robustness of MACT in terms of backbone models.
MACT is moderately efficient.
Table 5 compares MACT with other approaches in terms of the total number of prompting (generation) for each instance. For Binder and Dater, SC is performed a fixed number of times regardless of problem complexity. This results in a high number of prompting per instance, thus being inefficient. In contrast, MACT provides flexibility in generation, as the number of iterations depends on the problem’s complexity. For instance, most questions can be solved within three steps for WTQ (see our analysis in A.3), this results in a total of 25 times prompting3332 steps involve action and execution generation, with each five times, plus last step five times of action generation: 2*(5+5)+5. at most for each instance. If we incorporate the efficiency optimization module, which potentially saves up to one-third of the iterations (see Section 6), the total number of generations per instance is even lower (approximately 15), making MACT comparable to other approaches in terms of efficiency. The iterative nature of MACT can lead to a higher upper-bound of generation. However, it also allows for more complex problems, making the approach more tailored to real-life needs.
c Prompting number per instance
@@
Table 5: Prompting number of different approaches. We show lower and upper-bound if not deterministic.
6 Analysis
We conduct various analyses of our framework to back up our claims and contributions. Unless mentioned otherwise, all analyses are performed using Qwen-72b as $M_{p}$, CodeLlama-34b as $M_{c}$, the number of action generation $k=5$, and selection model $f_{s}=\texttt{SC}$. To explicitly analyze the effects of multi-agent collaboration with tool use, we do not use efficiency optimization, which means all instances undergo the iterative collaboration between $M_{p}$ and $M_{c}$ with tool use. Further analysis to support our choices of the sampling size $k$ and the maximum number of iterations $I$ are in A.2 and A.3.

### v2

@@ Abstract
Complex table question answering (TQA) aims to answer questions that require complex reasoning, such as multi-step or multi-category reasoning, over data represented in tabular form. Previous approaches demonstrate notable performance by leveraging either closed-source large language models (LLMs) or fine-tuned open-weight LLMs. However, fine-tuning LLMs requires high-quality training data, which is costly to obtain. The use of closed-source LLMs poses accessibility challenges and leads to reproducibility issues. In this paper, we propose Multi-Agent Collaboration with Tool use (MACT), a framework that requires neither fine-tuning nor closed-source models. In MACT, a planning agent and a coding agent that also make use of tools collaborate for TQA. MACT outperforms previous SoTA systems on three out of four benchmarks and performs comparably to the larger and more expensive closed-source model GPT-4 on two benchmarks, even when using only open-weight models without any fine-tuning. Our extensive analyses prove the effectiveness of MACT’s multi-agent collaboration in TQA. We release our code publicly. [^1] https://github.com/boschresearch/MACT
@@ 1 Introduction
The goal of table question answering (TQA) is to answer a question based on data represented in tabular form, optionally also using additional textual context. Recent studies on TQA focus more and more on complex instances, as they are ubiquitous in table data analysis (Zhu et al., 2021; Zhang et al., 2024b; Lu et al., 2023). Solving those complex instances requires performing multiple reasoning steps and/or employing different reasoning strategies (Ghosal et al., 2023). We refer to these aspects as multi-step and multi-category reasoning, respectively. An example requiring both types of reasoning is shown in the upper left part of Figure 1. To answer the question about the percentage change, a system first needs to use factual knowledge to extract countries in Europe. Then, numerical reasoning is applied to calculate the percentage change and to carry out the comparison.
@@
One popular approach for addressing those complex instances in TQA is planning, where step-wise plans are generated and used to guide the reasoning process (Zhang et al., 2024c; Wang et al., 2024; Wu and Feng, 2024; Zhu et al., 2024; Zhao et al., 2024). State-of-the-art works in this direction either fine-tune open-weight large language models (LLMs) Wu and Feng (2024); Zhu et al. (2024) or prompt closed-source commercial LLMs Wang et al. (2024); Zhang et al. (2024c); Zhao et al. (2024). However, fine-tuning requires high-quality data, which is usually expensive to obtain (Zhu et al., 2021). Prompting closed-source commercial LLMs can also be costly and poses challenges to reproducibility. To the best of our knowledge, existing methods leverage a single LLM to perform planning and reasoning, which is sub-optimal in particular if the LLM does not excel at mathematical reasoning or coding (Wu and Feng, 2024). These models struggle with answering questions requiring complex reasoning.
To address these challenges, we propose MACT, a multi-agent collaboration framework with tool use, which neither depends on closed-source LLMs nor requires fine-tuning. In fact, its backbone LLMs can be exchanged flexibly. It incorporates two agents (a planning agent and a coding agent) and a set of tools (a Python interpreter, a calculator and Wikipedia search). The planning agent performs online planning, i.e., it generates a plan iteratively. This breaks down complex problems and helps to address multi-step reasoning. The coding agent and the tool set assist with generating faithful intermediate results. The agents work in a collaborative setting, addressing the challenges of multi-category reasoning as all agents can concentrate on the reasoning types they excel in. An efficiency optimization module which allows the framework to take informed shortcuts.
We conduct experiments on four popular TQA benchmarks that include complex TQA instances. Our framework outperforms previous SoTA systems on three out of four benchmarks. It achieves comparable results to GPT-4 on two benchmarks even when using only open-weight models without any fine-tuning. In comparison to fine-tuned SoTA TQA systems, it demonstrates considerably better generalizability across datasets. Our analysis proves the effectiveness of our proposed collaborative setting of specialized agents. We find that the efficiency optimization module can save up to 33% of iterations without performance degradation.
@@ 2 Related Work
Table 1: Comparing MACT with previous works. NL plan stands for using natural language to encode a plan, cal=calculator, and Wiki=Wikipedia.
@@
We categorize previous work into three groups based on planning strategies: Heuristic coarse-grained planning consists of two pre-defined steps of retrieving and aggregating Ye et al. (2023); Zhou et al. (2024). Online global planning generates a plan in the first iteration and revises it in the next one Zhao et al. (2024). Online iterative planning conditions the generation of the next step on the executed results of previous steps Zhang et al. (2023a); Wang et al. (2024). We opt for online iterative planning for complex TQA, as it provides for more fine-grained steps during problem solving (in contrast to heuristic planning) and emphasizes the dependency among steps (in contrast to online global planning), which is crucial in complex TQA. In contrast to previous work using iterative planning, we introduce an efficiency optimization module to minimize the costs of the framework. To learn how to generate plans, previous work either depends on fine-tuning (Wu and Feng, 2024; Zhu et al., 2024), or strong closed-source models, combined with in-context learning (Wang et al., 2024; Zhang et al., 2023a; Zhao et al., 2024). By contrast, MACT generates effective plans using either closed-source or open-weight models, without the need for fine-tuning.
@@
In multi-agent collaboration settings, multiple AI entities collaborate towards a common goal (Talebirad and Nadiri, 2023). We use the term agent to refer to LLMs that interact with executable tools, following Qiao et al. (2024). As far as we know, none of the previous works in TQA utilize multi-agent collaboration. The approaches most closely related to ours explore collaboration among homogenous agents, i.e., all agents use the same backbone but are prompted differently (Liu et al., 2023; Zhao et al., 2024). The effectiveness of this approach relies heavily on strong (closed-source) backbone models. Our work explores multi-agent collaboration for TQA, without any constraint of model type.
@@ 3 Method
We propose MACT, a Multi-Agent Collaboration framework enriched with a set of Tools for TQA. Figure 1 provides an overview of the framework. It consists of four major modules: a memory $S$, a planning agent $M_{p}$, a coding agent $M_{c}$, and a tool set $T$. $M_{p}$ and $M_{c}$ are instantiated by (potentially) different LLMs. They collaborate through five core stages: action generation, action selection, tool selection/code creation, observation computation, and memory state update. The stages are executed iteratively for a maximum of $I$ iterations, where $I$ is a hyper-parameter. We control the overall efficiency of the collaboration via an efficiency optimization module. For a new TQA instance, we initialize the memory state $s_{0}=(\text{table},\text{question},\text{texts})$, i.e., with the input table, the question, and (if given) textual context. All parts of the memory are represented as strings. To represent the table as a string, we use pipes as column separators.
@@ 3.1 Action Generation
Retrieval textual description of what to retrieve, e.g., “sale numbers of 2019” $t=Python$ $M_{c}$ Python interpreter is run on generated code
@@
The intents Retrieval and Calculation are commonly seen in previous works (Gemmell and Dalton, 2023; Zhu et al., 2024). We use Retrieval for any operations extracting information from a table, including direct querying, filtering, and grouping. Instructions that require calculation, counting or comparison are captured by Calculation. To fulfill the possible need for external (factual) knowledge that is not present in the table or textual context, we add the intent Search, which performs Wikipedia searches to retrieve informative text passages. Read covers the need for contextual reasoning in table-text QA. It refers to instructions involving retrieving information from the texts provided as part of TQA instances. The intent Finish stops $M_{p}$ from generating more actions and ends the iterative execution of our framework, providing the final answer in the corresponding instruction. Lastly, we use an intent called Ask to retrieve an answer based on the internal knowledge of the planning agent. If $M_{p}$ fails to generate a valid action at an iteration, it will continue to generate the action at the next iteration until reaching the maximum iteration number $I$, and return the most common prediction directly from $M_{p}$ as the final answer (See section 3.7).
@@ 3.3 Tool Selection and Use
The tool needed for executing action $a^{*}_{i}$ depends on the intent of the action (see columns “Tool Selection” and “Tool Use/Execution” of Table 2). To address the intents Search, Calculation and Retrieval, we introduce a set of tools $T=\{T_{srh},T_{cal}\,T_{python}\}$. $T_{srh}$ is an API function for Wikipedia search (Shinn et al., 2023) from Langchain. [^2] https://rubydoc.info/gems/langchainrb/Langchain/Tool/Wikipedia The API takes a target entity specified in the instruction and returns the first paragraph of the corresponding Wikipedia entry. $T_{cal}$ is a calculator, powered by a Python interpreter. It takes a formula generated by $M_{p}$ and outputs the answer. Note that the instruction of Calculation can also be a textual description, such as “Compute the average number of medals for each country in the table.” To better address these instructions, we introduce a coding agent $M_{c}$ with a Python interpreter, denoted as $T_{python}$ (see 3.4). We do not distinguish formulas and text description. This means any instructions with the intent Calculation will be firstly passed to $T_{cal}$. If $T_{cal}$ fails to execute the instruction, $T_{python}$ is applied.
The intent Retrieval is also addressed by $M_{c}$ and $T_{python}$, i.e., $M_{c}$ generates Python code based on a given instruction to retrieve target cells in the tables and the Python interpreter returns the executed results. Lastly, For Read, Ask and Finish, no tool is used, denoted as $t=Null$ in Table 2. The answers to the intents Read and Ask are queried via $M_{p}$. For Read, $M_{p}$ reads the answer from a given textual context. For Ask, it responds based on its internal knowledge. No execution is performed for Finish, as the intent ends MACT with a final answer in its instruction.
@@ 4 Experiments
We categorize the recent TQA systems into two groups indicating if they require LLM fine-tuning or not. We include the following baselines that require fine-tuning: OmniTab (Jiang et al., 2022), TableLlama Zhang et al. (2024a), ProTrix (Wu and Feng, 2024), TAT-LLM (Zhu et al., 2024) and TableLLM Zhang et al. (2024b). Except for OmniTab, which is backboned by BART (Lewis et al., 2020), all others build on top of LLaMA 7b (Touvron et al., 2023). Our set of TQA baselines that do not require fine-tuning includes Dater (Ye et al., 2023), Binder (Cheng et al., 2023), Chain-of-Table (Wang et al., 2024), ReAcTable (Zhang et al., 2024c), TabSQLify (Nahid and Rafiei, 2024), Plan-then-Reason (Wu and Feng, 2024), Mix-SC (Liu et al., 2023) and ARC Zhang et al. (2023b). They all rely on GPT-3.5-turbo.
@@
In MACT, the choice of the planning and coding agent is flexible as no fine-tuning is involved. We experiment with the best open-weight LLMs available at the time of writing: Qwen-2 72B (Yang et al., 2024) (planning agent) and CodeLLaMA-34B (coding agent). They are run on int4 and full precision, respectively. We further use GPT-3.5-turbo as both the planning and coding agents when comparing our method with other TQA systems that use this model. $\tau_{p}$ and $\tau_{c}$ are set to 0.6 for non-repetitive action generation. We set the action and code generation size $k$ to 5, following Liu et al. (2023). The maximum number of iteration $I$ is set to 7, based on empirical results on the development sets. For the efficiency component, we set $\alpha$ to 1 to ensure high confidence of the model. We explore the effects of different values of $\alpha$ in Section 6. Two NVIDIA A100 GPUs are used for running MACT.
@@ 5 Results
We first evaluate the performance of MACT in direct comparison to recent TQA systems using closed-source LLMs. Second, we examine how our method performs compared to fine-tuned open-weight LLMs. [^3] We did not run ARC as no code is available. Results for TAT-LLM is only reported on TAT as the model is specially designed for TAT dataset that features both tables and texts as inputs. Following prior work, we use exact match (EM) as the evaluation measure. Lastly, we discuss the efficiency of MACT.
@@
Table 3: Exact Match results of models without fine-tuning. SCT refers to SCITAB. The models are grouped by the LLM they use as their backbone (top: GPT-3.5; middle: open-weight LLMs as indicated in parentheses). Performances marked with * are taken from the original paper. Performances marked with †are taken from Wu and Feng (2024), Zhu et al. (2024), Zhang et al. (2023b) and Lu et al. (2023) for each dataset. We bold the best performances in each group.
@@
The middle part of Table 3 provides the results of MACT using specialized agents (MACT (Qw+CL): Qwen-2 as planning agent, CodeLLaMA as coding agent). As baselines, we compare a setting without a specialized coding agent (MACT (Qw+Qw)) and a setting without a general planning agent (MACT(CL+CL)). As further baselines, we use the two LLMs on their own as well in combination, with Chain-of-thought prompting (Wei et al., 2022). For combination, the single models are prompted five times and combined using SC, as in Liu et al. (2023). This is a direct multi-agent baseline without collaboration and tool use. Both SC(Qw+CL) and MACT (Qw+CL) achieve higher EM scores than individually prompting Qwen and CodeLLaMA, demonstrating the positive effect of using multiple agents for planning and coding. Importantly, MACT (Qw+CL) outperforms SC(Qw+CL) by approximately 6 EM points on average across all datasets, highlighting the superiority of our collaboration technique over simply taking the most frequent predictions from two independent agents. We also find that having an expert coding agent for code generation (MACT (Qw+Qw) vs. MACT (Qw+CL)) improves performance considerably.
MACT with open-weight models delivers comparable performance as closed-source systems.
As shown in Table 3, by comparing MACT (Qw+CL) with TQA systems that rely on closed-source LLMs (in the upper part of Table 3), we find that our model outperforms the examined TQA systems for three out of four datasets. Our multi-agent TQA system is more cost-efficient and straightforward to replicate, while delivering superior performance compared to closed-source TQA models. In addition, we show two upper-bounds in the bottom part of Table 3: The performance of human annotators and directly prompting GPT-4 with the table and question. As shown in Table 3, GPT-4 has an advantage on TAT and SCITAB. On WTQ, we observe comparable performances between MACT (QW+CL) and GPT-4. On CRT, our method even outperforms GPT-4 by 5.7%. CRT is the most complex dataset, requiring multi-step and multi-category reasoning, which direct inference with GPT-4 cannot generally solve. Our step-wise collaborative planning setting is well-suited to such settings. In contrast, there is a large gap between MACT and human performance in SCITAB. SCITAB collects data from scientific papers, in which abbreviations and domain-specific terms are common. These can pose challenges to current systems and models. In TAT, MACT often finds the correct answer but struggles to output it in the correct format (see Sec. 6).
Table 4 compares our framework with prior fine-tuned TQA models. We present results for MACT with different planning and coding agents: Our standard setting (LlaMA-7b+CS-7b) as well as with a stronger planner (Qw-7b+CS-7b). In general, for fine-tuned TQA models, their performance on the dataset used for fine-tuning is rather high while they suffer from a considerable drop in EM when tested on other datasets. This observation is in line with the findings by Zhang et al. (2024a) and Huang et al. (2024). In contrast, MACT does not use fine-tuned models and can, thus, be applied to any dataset with a good generalization performance. MACT demonstrates comparable results to Protrix when using LlaMA-7b as the planning agent, though it has not been fine-tuned. As expected, using a better planning agent leads to better results. This also shows the robustness of MACT in terms of backbone models.
@@
TableLlama (LlaMA-7b) c 29.9 17.4 26.9 38.6
Protrix (LlaMA-7b) c 48.9* 26.8 40.2 42.4
TAT-LLM (LlaMA-7b) c - 69.6* - -
@@
Table 4: Exact Match Results of MACT using different LLM agents in comparison to fine-tuned TQA models. Performances marked with * refer to the in-domain setting (where fine-tuning took place). SCT refers to SCITAB. CS refers to the deepseek-coder model.
MACT adapts computational cost to instance complexity.
Table 5 compares MACT with other approaches in terms of the total number of LLM calls for each instance. For Binder and Dater, SC is performed a fixed number of times regardless of problem complexity. This results in a high number of LLM calls per instance, making them inefficient. In contrast, MACT provides flexibility in generation, as the number of iterations depends on the problem’s complexity. For instance, most questions can be solved within three steps for WTQ (see our analysis in A.3). This results in a total of at most 25 LLM calls 4442 steps involve action and execution generation, with each five times, plus last step five times of action generation: 2*(5+5)+5. for each instance. If we incorporate the efficiency optimization module, which potentially saves up to one-third of the iterations (see Section 6), the total number of LLM calls per instance is even lower (approximately 15), making MACT comparable to other approaches in terms of efficiency. The iterative nature of MACT can lead to a higher upper-bound of LLM calls. However, it also allows for solving more complex problems, making the approach more tailored to real-life requirements.
c Number of LLM calls per instance
@@
Table 5: Number of LLM calls for different approaches. We show lower and upper-bounds if not deterministic.
@@
6 Analysis
We conduct various analyses of our framework to back up our claims and contributions. Unless mentioned otherwise, all analyses are performed using Qwen-2 72B as $M_{p}$, CodeLlama-34B as $M_{c}$, the number of action generation $k=5$, and selection model $f_{s}=\texttt{SC}$. To explicitly analyze the effects of multi-agent collaboration with tool use, we do not use efficiency optimization, which means all instances undergo the iterative collaboration between $M_{p}$ and $M_{c}$ with tool use. Further analysis to support our choices of the sampling size $k$ and the maximum number of iterations $I$ are in A.2 and A.3. A case study can be found in 5.
Figure 2: Distribution of action intents by dataset.